OpenStudy (anonymous):

Find the equation of the parabola with focus (2, 3) and directrix y = -1.

OpenStudy (anonymous):

@Hero

hero (hero):

Have you not figured out how to do these yet?

OpenStudy (anonymous):

I asked you for clarification on the one you said I was doing wrong & you didnt reply.

hero (hero):

I was about to reply but I discovered that satellite was helping you.

OpenStudy (anonymous):

The way he is helping is by the graphs and I don't understand that way . I prefer trying to square it I just don't know how to .

hero (hero):

So you don't know how to muliply this: (y - 2)(y - 2)

hero (hero):

That's the same as \((y - 2)^2\)

hero (hero):

Working on these assumes you know how to multiply binomials.

OpenStudy (anonymous):

yes I know how to multiply that .

hero (hero):

Well, when you multiply two binomials that are the same, you are squaring them.

hero (hero):

I hope that helps you.

OpenStudy (anonymous):

oh okay soo .. this problem would begin by doing (x-2)^2 (x-3)^2 ?

hero (hero):

If you're going to do it, you might as well be thorough and write it all out. Otherwise, I'm going to assume that you are making a mistake.

hero (hero):

x-2)^2 + (x-3)^2 = (x - x)^2 + (y -(-1))^2

OpenStudy (anonymous):

x^2 - 4 + x^2-9 = x^2 +y^2 +1 ???

hero (hero):

squaring binomials is not the same as a difference of squares: \((x - 2)^2 \ne x^2 - 4\)

OpenStudy (anonymous):

Im so sorry , but I am confused !

hero (hero):

\((x - 2)^2 = (x - 2)(x - 2)\) \(x^2 - 4 = (x + 2)(x - 2)\)

hero (hero):

Hopefully, that will clear up some of the confusion

hero (hero):

\((x - 3)^2 = (x - 3)(x - 3)\) \(x^2 - 9 = (x + 3)(x - 3)\)

OpenStudy (anonymous):

I simplified your original problem too : (x-3)^2 + (x-2)^2 = (y+1)^2 does that help anything ?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, but go ahead and expand it from there

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

like just on the first binomial ? so i know how to do it ..

hero (hero):

I thought you said you know how to multiply binomials. If you are confused just write it out Do you know how to multiply (y - 2)(y - 2)?

OpenStudy (anonymous):

y^2 +4 ?

hero (hero):

apply FOIL to avoid confusion

OpenStudy (anonymous):

So that is : Y(y) + 2(y) .. ? 2(2)

OpenStudy (anonymous):

I am missing one .. can you help me ?

hero (hero):

@Faith_Rochelle, if you don't know how to multiply binomials, then you shouldn't be doing these.

OpenStudy (anonymous):

@Hero the home school I do my mom is supposed to be my teacher and she dont know how to do these. can you please just explain it ?

hero (hero):